Find the area of the triangle determined by the points (2,-1,0),(2,0,4) and (0,2,3)


1
Expert's answer
2021-08-12T15:26:03-0400

Let "\overline {u}=(2-2,0-(-1),4-0)=(0,1,4)" be the vector joining "(2,-1,0)" and "(2,0,4)." Let "\overline {v}=(0-2,2-(-1),3-0)=(-2,3,3)" be the vector joining "(2,-1,0)" and "(0,2,3)."

Let us find the vector product


"\overline{u}\times\overline{v}=
\begin{vmatrix}i & j & k\\
0 & 1 & 4\\
-2 & 3 & 3
\end{vmatrix}
=-9i-8j+2k."

It follows that the area of the triangle determined by the points"(2,-1,0),(2,0,4)" and "(0,2,3)" is equal to


"A=\frac{1}{2}|\overline{u}\times\overline{v}|=\frac{1}{2}\sqrt{(-9)^2+(-8)^2+2^2}=\frac{1}{2}\sqrt{149}\approx 6.1" (square units)